# EByFTVeS: Efficient Byzantine Fault Tolerant-based Verifiable Secret-sharing in Distributed Privacy-preserving Machine Learning

### v1

@@ Abstract
Verifiable Secret Sharing (VSS) has been widespread in Distributed Privacy-preserving Machine Learning (DPML), because invalid shares from malicious dealers or participants can be recognized by verifying the commitment of the received shares for honest participants. However, the consistency and the computation and communitation burden of the VSS-based DPML schemes are still two serious challenges. Although Byzantine Fault Tolerance (BFT) system has been brought to guarantee the consistency and improve the efficiency of the existing VSS-based DPML schemes recently, we explore an Adaptive Share Delay Provision (ASDP) strategy, and launch an ASDP-based Customized Model Poisoning Attack (ACuMPA) for certain participants in this paper. We theoretically analyzed why the ASDP strategy and the ACuMPA algorithm works to the existing schemes. Next, we propose an [E]fficient [By]zantine [F]ault [T]olerant-based [Ve]rifiable [S]ecret-sharing (EByFTVeS) scheme. Finally, the validity, liveness, consistency and privacy of the EByFTVeS scheme are theoretically analyzed, while the efficiency of the EByFTVeS scheme outperforms that of the-state-of-art VSS scheme according to comparative experiment results.
Index Terms:
Byzantine Fault Tolerant, Model Poisoning, Consistency, Verifiable Secret Sharing, Distributed Privacy-preserving Machine Learning.
@@ I Introduction
Verifiable Secret Sharing (VSS)[vss-original] is a critical data privacy-preserving technique that extends the concept of secret sharing by ensuring the correctness and integrity of the distributed secret shares. In a standard VSS scheme, a dealer divides her sensitive data into multiple shares and distributes them among a group of participants. The sensitive data can only be reconstructed after a sufficient number of shares are combined. By incorporating commitment mechanism, VSS allows each participant to independently verify the validity of the received share, ensuring that malicious dealers and participants are negligible to provide invalid shares. The VSS is a paramount technique in building robust and secure distributed systems, such as secure multiparty computation and threshold cryptography, as it prevents a malicious dealer from sabotaging the secret-sharing process and ensures that the sensitive data can be reliably reconstructed by any qualified subset of honest participants.
The property of verifiability ensures the integrity of distributed shares even in the presence of malicious dealers and participants, holds significant promise in Distributed Privacy-preserving Machine Learning (DPML) applications. For an instance, a verifiable and private patient record sharing scheme was proposed in a crowdsourcing-based healthcare system [ss-healthcare], highlighting the requirement of data integrity of patient medical records, while a collusion-deterrent verifiable secret sharing mechanism was proposed to discourage parties from colluding and recovering the dealer’s secret [eurocrypt-vss].
In fact, two significant hurdles in the widespread adoption of VSS lies in its inherent inconsistency of the commitment, and computational and communicational overhead. On one hand, since each participant independently verifies the validity of the received commitments with the shares, malicious dealers can easily provide different commitments to different participants, while malicious participants can replay historical shares and commitments to other participants when the secret requires to be reconstructed. In other words, the consistency of both commitments and shares cannot be guaranteed among the participants. The inconsistent shares can lead the model poisoning attack which makes the distributed machine learning model difficult to be converged. On the other hand, the cryptographic primitives implemented to ensure the verifiability of shares like polynomial commitments, usually involve complex mathematical operations of abstract algebra operations. These computations can be resource intensive, demanding considerable processing power from both the dealers during in the sharing phase and the participants during verification and reconstruction phase. In addition, broadcasting the share-validity evidences, along with the shares themselves, can lead to substantial communication costs, especially in scenarios with a large number of participants or when dealing with large shares. These consistent, computational and communicational demands can pose practical limitations, particularly in DPML applications that usually require high throughput to support huge size of model parameters [mp-spdz].
In recent years, Byzantine Fault Tolerance (BFT) has been gradually involved in promising the consistency of commitments and accelerating VSS-based secure multiparty computation schemes. An efficient verifiable secret sharing scheme [basuEfficientVerifiableSecret2019] was proposed to ask dealers and participants to broadcast shares and commitments onto a BFT system. Since each dealer can only provide her share and commitment once in a training round of the DPML application on the BFT system, a malicious dealer or participant is difficult to provide inconsistent or out-of-date shares and commitments anymore. Unfortunately, we explore a new adaptive share delay provision strategy which enables malicious dealers to compute customized shares and postpone to broadcast these ’meticulous constructed’ shares. Malicious dealers can still apply this strategy to launch model poisoning attack, while victim participants are proven to imperceptibly pass the verification for these customized shares.
The contribution of this paper is summarized as follows.
  • We first explore an Adaptive Share Delay Provision (ASDP) strategy. A malicious dealer can generate any shares as her wish to victim participants by observing the on-chain shares at the start. All of the victim participants are proven to accept the malicious dealer’s shares in theory.
  • Next, we launch a [A]SDP-based [Cu]stomized [M]odel [P]oisoning [A]ttack (ACuMPA) for certain victim participants using ASDP strategy. We formally analyzed that the ACuMPA algorithm can pass the typical model poisoning defense mechanism such as cosine similarity check.
  • Finally, we propose an [E]fficient [By]zantine [F]ault [T]olerant-based [Ve]rifiable [S]ecret-sharing (EByFTVeS) scheme. This scheme forces participants to only aggregate those shares whose commitments are provided in the BFT system at a special time period. This restriction makes the ACuMPA not valid anymore.
  • We formally analyze the validity, liveness, consistency, privacy of the EByFTVeS scheme, while experimental analysis shows that the efficiency of EByFTVeS outperforms that of the state-of-the-art VSS-based distributed privacy-preserving machine learning schemes.
The remainder of this paper is organized as follows. Section II recalls the related work. In Section III, we present the preliminaries. Then the models and goals are both formulated in Section IV. Section V and Section VI propose the SDP strategy and the ACuMPA scheme, as well as explaining why the strategy and attack are valid in theory. In Section VII, the EByFTVeS scheme is proposed and formally analyzed. Section VIII demonstrates the efficiency of the EByFTVeS scheme. Section IX draws the conclusion.
@@ II Related Works
Verifiable Secret Sharing (VSS) plays a fundamental role in constructing robust Secure Multiparty Computation (SMC) protocols, particularly when dealing with active adversaries who may deviate arbitrarily from the protocol [vss-original, vss-feldman, vss-benaloh]. By ensuring that the initial secret-sharing phase is verifiable, VSS prevents a malicious party from distributing inconsistent or invalid shares that could compromise the entire computation.
Related works in this area focus on designing efficient VSS schemes that can be seamlessly integrated into various SMC frameworks, minimizing the overhead while maximizing security guarantees against a variety of adversarial models. For instance, BEAR [ssfl-jsac] pioneers the use of VSS to securely compute gradient distances for KRUM aggregation, leveraging polynomial operations to ensure robustness. Similarly, ELAS [elas-beaver] incorporates secret sharing combined with Beaver triples to calculate the $l_{2}$ norm for achieving secure aggregation. Furthermore, RFLPA [RPLPA-nips] introduces a novel dot-product-based aggregation rule, leveraging secret sharing to compute cosine similarity and effectively defend against poisoning attacks, showcasing the versatility of VSS in distributed machine learning scenarios. Frankly, almost all of the aforementioned works have to apply Beaver triples to support multiplication operations. Since the construction of Beaver triples relies on multiple rounds of oblivious transfer, which is heavy to compute and communicate intermediate parameters, therefore, the efficiency of these works has yet to be a great challenge.
@@
Related works in this domain concentrate on developing efficient and secure threshold signature schemes, threshold encryption schemes, and distributed key generation protocols that rely on VSS as a fundamental building block. Li et al.[DKG-VSS] proposed a distributed key generation protocol based on VSS, which reduces the communication overhead during the key generation process. Similarly, TLARDA[TLARDA-audit] introduced a remote data integrity auditing framework in decentralized environments by leveraging two VSS schemes, achieving a significant reduction in communication complexity from $\mathcal{O}(n)$ to $\mathcal{O}(1)$. Ascina [ascina] utilized VSS to offload tag computation tasks in industrial cloud storage systems, effectively lowering computational overhead. While these studies have made strides in addressing challenges such as reducing communication and computation costs, ensuring robustness against malicious participants, and enhancing the efficiency of the VSS phase, they still rely heavily on the verification of shares to maintain the consistency of secrets. As a result, there remains significant potential for further improving the overall efficiency of these approaches.
@@
Related research explores how VSS can be adapted and optimized for use in various BFT consensus algorithms, addressing challenges like maintaining consistency in the presence of malicious dealers, ensuring validity, privacy and liveness properties, and minimizing the performance impact of the VSS component on the overall system’s efficiency and fault tolerance capabilities. Although a BFT-based VSS scheme [basuEfficientVerifiableSecret2019] was recently proposed to force dealers to broadcast the commitments before shares onto a BFT system for maintaining the consistency of the shares. This scheme was still hard to resist against model poisoning attacks based on our exploration that will be shown in Section V.
@@ III-A VSS
VSS [shamir-secret-sharing] comprises of two kinds, the Shamir VSS and additive VSS. In the following sections, we mainly use Shamir VSS as examples, though our scheme is also valid for additive VSS. A typical Shamir secret sharing has two parameters, usually named $f$ and $n$, where $f$ denotes the minimal size of a subset required to recover secret, and $n$ presents the total number of participants. When $f=n$, the function of Shamir VSS can be viewed same as that of additive VSS.
@@ III-B Byzantine Fault Tolerance (BFT) System
Byzantine Fault Tolerance (BFT) addresses the classic Byzantine Generals Problem [lamport2019byzantine], where a group of generals must reach consensus on a battle plan despite the presence of traitors who may send conflicting messages. In BFT systems, this translates to achieving agreement among participants when some might be faulty or malicious, potentially sending inconsistent information to different parts of the network.
Classical BFT systems are categorized based on their network assumptions: synchronous [chan2018pili, abraham2020sync], asynchronous [duan2023fin, gao2022dumbo], and partially synchronous. Partially synchronous [pbft, liu2025group, hca] models offer the most practical balance between theoretical guarantees and real-world applicability. We select the Practical Byzantine Fault Tolerance (PBFT) [pbft] algorithm in this model as the foundation for the design of our consensus-based local gradient exchange algorithm, owing to its fundamental and widely applicable nature. PBFT follows a leader-based consensus paradigm [liu2025abse] and can tolerate up to $f$ Byzantine faults in a system with at least 3ff+1 participants, requiring a quorum of at least $2f+1$ of the participants to agree.
@@ III-C The VSS-based Distributed Privacy-preserving Machine Learning
Distributed privacy-preserving machine learning collaborates multi-participants to train a machine learning model together. It is usually used when the data of each participants are not sufficient enough to make the model convergence. In order to preserve the privacy of the trained parameters, VSS is usually used when participants aggregate all the local trained parameters from her own and others [ssfl-jsac]. The specification of the VSS-based distributed privacy-preserving machine learning algorithm is listed as below.
@@ V Adaptive Share Delay Provision
Since each participant has to independently aggregate $th$ shares by running the function of MPA in algorithm 1 for training the model, malicious dealers can meticulously elaborate their model parameters to resist against the common model poisoning prevention mechanisms, like cosine similarity-based [fltrust] or $L$ regularization-based [krum] gradient aggregation.
@@ VII The [E]fficient [By]zantine [F]ault [T]orlerant-based [Ve]rifiable [S]ecret-sharing (EByFTVeS) Scheme
In this section, we apply a four-phase consensus algorithm based on Practical Byzantine Fault Tolerance (PBFT) [pbft] to propose the EByFTVeS scheme against model poisoning attack. We then analyze the validity, liveness, consistency, and privacy properties of the EByFTVeS scheme.
@@ VII-A The EByFTVeS Algorithm
PBFT [pbft] was chosen as it is a fundamental and widely applicable BFT protocol. However, we note that this is not the only feasible option, and other BFT protocols [dolphin, hotstuff] could also be used. In the normal-case operation of the classical PBFT, participants directly submit requests to a primary participant, a process that is not feasible in the context of the DPML. To address this, we introduce a modification by adding an additional Pre-Propose phase before the conventional three phases of Pre-Prepare, Prepare, and Commit. Note we only sketch the major workflow for normal-case operation here; mechanisms such as checkpointing are excluded on account of their alignment with the original protocol (the array we created should also be considered during garbage collection).
@@ VII-B Theoretical Analysis
We consider both intra-view and inter-view scenarios. Case 1 (Intra-view Validity:): We prove this by contradiction. Suppose honest participants $p_{i}$ and $p_{j}$ commit different proposals propprop and $prop^{\prime}$ respectively with the same sequence number $sq$ in view $v$. For participant $p_{i}$ to commit propprop, it must have received $2f+1$ matching Commit messages for some digest $\delta=h(prop)$ (i.e., $committed\text{-}local(\delta,v,sq)=true$). Similarly, for participant $p_{j}$ to commit $prop^{\prime}$, it must have received $2f+1$ matching Commit messages for some digest $\delta^{\prime}=h(prop^{\prime})$ (i.e., $committed\text{-}local(\delta^{\prime},v,sq)=true$). Since there are at most $f$ Byzantine participants, By the quorum intersection property (2(2f+1)n=(4f+2)(3f+1)=f+12(2f+1)-n=(4f+2)-(3f+1)=f+1), there exists at least one honest participant $p^{*}$ that sent both Commit messages for $\delta$ and $\delta^{\prime}$. However, honest participants only send one Commit message per sequence number per view, leading to a contradiction.
@@ VIII-A1 Experimental Settings
Datasets and Models. We evaluate the performance of the proposed ACuMPA scheme on three widely used visual recognition benchmark datasets: MNIST [lecun1998mnist], Fashion-MNIST [xiao2017fashion], and CIFAR-10 [krizhevsky2009learning].
@@
For model evaluation, we employ three popular neural network architectures: a Convolutional Neural Network (CNN) [chua1997cnn], ResNet [he2016deep], and AlexNet [krizhevsky2012imagenet]. Detailed specifications of the model structures are presented in Table II.

### v2

@@ Abstract
In the realm of Distributed Privacy-preserving Machine Learning (DPML), Verifiable Secret Sharing (VSS) serves as a fundamental primitive for guaranteeing data integrity against malicious actors. Despite its utility, VSS-based DPML frameworks frequently struggle with issues related to view consistency and prohibitive resource consumption. While recent integrations of Byzantine Fault Tolerance (BFT) protocols have attempted to mitigate these consistency issues, we identify a residual vulnerability inherent in current designs. We demonstrate that adversaries can exploit the lack of strict synchronization in share provision through a strategy we term Adaptive Share Delay Provision (ASDP). This strategy enables a novel adversarial technique, the ASDP-based Customized Model Poisoning Attack (ACuMPA), which allows attackers to tailor malicious gradients for specific targets. We provide a rigorous theoretical foundation explaining the efficacy of ACuMPA against existing safeguards. To neutralize this threat, we introduce EByFTVeS, a robust framework that couples BFT consensus with verifiable secret sharing. By enforcing a strict temporal lock on share commitments via the consensus mechanism, EByFTVeS effectively precludes the delayed injection of poisoned shares. Our theoretical analysis confirms the protocol’s properties regarding validity, liveness, consistency, and privacy. Furthermore, comprehensive empirical evaluations indicate that EByFTVeS achieves superior computational efficiency compared to state-of-the-art VSS alternatives.
@@ I Introduction
Secure Multiparty Computation (SMC) and threshold cryptography rely heavily on Verifiable Secret Sharing (VSS) [1] to establish trust in distributed environments. By extending classical secret sharing with cryptographic commitments, VSS empowers participants to verify the correctness of received shares without revealing the underlying secret. This capability is indispensable for robust distributed systems, ensuring that a dealer cannot distribute invalid or conflicting shares to honest participants. Consequently, VSS has emerged as a cornerstone technology for ensuring data integrity and preventing sabotage in collaborative computational tasks.
The application of VSS within Distributed Privacy-preserving Machine Learning (DPML) has garnered significant attention due to its potential to secure sensitive data updates, such as gradient exchanges. For instance, recent works have leveraged VSS to secure patient records in crowdsourcing healthcare systems [2] and to deter collusion among computing parties [3]. However, the practical deployment of these schemes is often hindered by two pervasive challenges: the equivocation of commitments and the exorbitant overhead associated with cryptographic operations.
First, the issue of consistency remains a critical vulnerability. In the absence of a unified broadcast channel, a malicious dealer acts as a single point of failure by potentially sending conflicting commitments to different participants. Similarly, compromised participants may attempt to replay obsolete shares during the reconstruction phase. Such inconsistencies can severely disrupt the learning process, facilitating model poisoning attacks that prevent the global model from converging. Second, the computational complexity of VSS primitives—particularly those involving homomorphic polynomial commitments—imposes a severe performance penalty. The necessity of broadcasting validity proofs alongside raw shares exacerbates communication latency, rendering standard VSS protocols ill-suited for high-throughput DPML scenarios involving massive parameter sets [4]. While some benchmarking frameworks have attempted to alleviate online costs by offloading Beaver triple generation to an offline phase, the online verification of secret consistency remains a bottleneck.
To address the consistency dilemma, researchers have begun incorporating Byzantine Fault Tolerance (BFT) mechanisms into VSS workflows [5]. By mandating that dealers publish shares and commitments via a BFT-ordered ledger, these protocols aim to create an immutable record of the sharing phase. Theoretically, this prevents a dealer from equivocating or updating shares mid-round. However, we reveal that these measures are insufficient against a sophisticated adversary. We uncover a timing-based vulnerability: the Adaptive Share Delay Provision (ASDP). By strategically delaying the broadcast of shares until honest participants have revealed their contributions, a malicious dealer can craft and inject ”customized” shares that bypass validity checks while poisoning the model.
In this work, we present a comprehensive solution to these challenges. Our contributions are fourfold:
  • We formally define the Adaptive Share Delay Provision (ASDP) strategy. We demonstrate how a malicious dealer can leverage information asymmetry to generate arbitrary shares for victim participants, and we prove that these fabricated shares will inevitably be accepted by the victims under current protocol definitions.
  • We formulate the ASDP-based Customized Model Poisoning Attack (ACuMPA). By utilizing the ASDP strategy, we show how an adversary can circumvent standard defenses, such as cosine similarity filtering, to corrupt the global model aggregation process.
  • We propose EByFTVeS, an optimized VSS scheme integrated with BFT consensus. The core innovation of EByFTVeS lies in its consensus-driven synchronization, which mandates that only shares committed within a verified consensus window are eligible for aggregation. This mechanism effectively eliminates the time window required for the ACuMPA attack.
  • We provide a rigorous theoretical analysis of the proposed scheme, covering validity, liveness, consistency, and privacy. Empirical comparisons with existing state-of-the-art solutions confirm that EByFTVeS delivers significant improvements in both security and execution efficiency.
@@ II Related Works
Verifiable Secret Sharing (VSS) plays a fundamental role in constructing robust Secure Multiparty Computation (SMC) protocols, particularly when dealing with active adversaries who may deviate arbitrarily from the protocol [1, 6, 7]. By ensuring that the initial secret-sharing phase is verifiable, VSS prevents a malicious party from distributing inconsistent or invalid shares that could compromise the entire computation.
Related works in this area focus on designing efficient VSS schemes that can be seamlessly integrated into various SMC frameworks, minimizing the overhead while maximizing security guarantees against a variety of adversarial models. For instance, BEAR [8] pioneers the use of VSS to securely compute gradient distances for KRUM aggregation, leveraging polynomial operations to ensure robustness. Similarly, ELAS [9] incorporates secret sharing combined with Beaver triples to calculate the $l_{2}$ norm for achieving secure aggregation. Furthermore, RFLPA [10] introduces a novel dot-product-based aggregation rule, leveraging secret sharing to compute cosine similarity and effectively defend against poisoning attacks, showcasing the versatility of VSS in distributed machine learning scenarios. Frankly, almost all of the aforementioned works have to apply Beaver triples to support multiplication operations. Since the construction of Beaver triples relies on multiple rounds of oblivious transfer, which is heavy to compute and communicate intermediate parameters, therefore, the efficiency of these works has yet to be a great challenge.
@@
Related works in this domain concentrate on developing efficient and secure threshold signature schemes, threshold encryption schemes, and distributed key generation protocols that rely on VSS as a fundamental building block. Li et al.[11] proposed a distributed key generation protocol based on VSS, which reduces the communication overhead during the key generation process. Similarly, TLARDA[12] introduced a remote data integrity auditing framework in decentralized environments by leveraging two VSS schemes, achieving a significant reduction in communication complexity from $\mathcal{O}(n)$ to $\mathcal{O}(1)$. Ascina [13] utilized VSS to offload tag computation tasks in industrial cloud storage systems, effectively lowering computational overhead. While these studies have made strides in addressing challenges such as reducing communication and computation costs, ensuring robustness against malicious participants, and enhancing the efficiency of the VSS phase, they still rely heavily on the verification of shares to maintain the consistency of secrets. As a result, there remains significant potential for further improving the overall efficiency of these approaches.
@@
Related research explores how VSS can be adapted and optimized for use in various BFT consensus algorithms, addressing challenges like maintaining consistency in the presence of malicious dealers, ensuring validity, privacy and liveness properties, and minimizing the performance impact of the VSS component on the overall system’s efficiency and fault tolerance capabilities. Although a BFT-based VSS scheme [5] was recently proposed to force dealers to broadcast the commitments before shares onto a BFT system for maintaining the consistency of the shares. This scheme was still hard to resist against model poisoning attacks based on our exploration that will be shown in Section V.
@@ III-A VSS
VSS [14] comprises of two kinds, the Shamir VSS and additive VSS. In the following sections, we mainly use Shamir VSS as examples, though our scheme is also valid for additive VSS. A typical Shamir secret sharing has two parameters, usually named $f$ and $n$, where $f$ denotes the minimal size of a subset required to recover secret, and $n$ presents the total number of participants. When $f=n$, the function of Shamir VSS can be viewed same as that of additive VSS.
@@ III-B Byzantine Fault Tolerance (BFT) System
Byzantine Fault Tolerance (BFT) addresses the classic Byzantine Generals Problem [15], where a group of generals must reach consensus on a battle plan despite the presence of traitors who may send conflicting messages. In BFT systems, this translates to achieving agreement among participants when some might be faulty or malicious, potentially sending inconsistent information to different parts of the network.
Classical BFT systems are categorized based on their network assumptions: synchronous [16, 17], asynchronous [18, 19], and partially synchronous. Partially synchronous [20, 21, 22] models offer the most practical balance between theoretical guarantees and real-world applicability. We select the Practical Byzantine Fault Tolerance (PBFT) [20] algorithm in this model as the foundation for the design of our consensus-based local gradient exchange algorithm, owing to its fundamental and widely applicable nature. PBFT follows a leader-based consensus paradigm [23] and can tolerate up to $f$ Byzantine faults in a system with at least 3ff+1 participants, requiring a quorum of at least $2f+1$ of the participants to agree.
@@ III-C The VSS-based Distributed Privacy-preserving Machine Learning
Distributed privacy-preserving machine learning collaborates multi-participants to train a machine learning model together. It is usually used when the data of each participants are not sufficient enough to make the model convergence. In order to preserve the privacy of the trained parameters, VSS is usually used when participants aggregate all the local trained parameters from her own and others [8]. The specification of the VSS-based distributed privacy-preserving machine learning algorithm is listed as below.
@@ V Adaptive Share Delay Provision
Since each participant has to independently aggregate $th$ shares by running the function of MPA in algorithm 1 for training the model, malicious dealers can meticulously elaborate their model parameters to resist against the common model poisoning prevention mechanisms, like cosine similarity-based [24] or $L$ regularization-based [25] gradient aggregation.
@@ VII The [E]fficient [By]zantine [F]ault [T]orlerant-based [Ve]rifiable [S]ecret-sharing (EByFTVeS) Scheme
In this section, we apply a four-phase consensus algorithm based on Practical Byzantine Fault Tolerance (PBFT) [20] to propose the EByFTVeS scheme against model poisoning attack. We then analyze the validity, liveness, consistency, and privacy properties of the EByFTVeS scheme.
@@ VII-A The EByFTVeS Algorithm
PBFT [20] was chosen as it is a fundamental and widely applicable BFT protocol. However, we note that this is not the only feasible option, and other BFT protocols [26, 27] could also be used. In the normal-case operation of the classical PBFT, participants directly submit requests to a primary participant, a process that is not feasible in the context of the DPML. To address this, we introduce a modification by adding an additional Pre-Propose phase before the conventional three phases of Pre-Prepare, Prepare, and Commit. Note we only sketch the major workflow for normal-case operation here; mechanisms such as checkpointing are excluded on account of their alignment with the original protocol (the array we created should also be considered during garbage collection).
@@ VII-B Theoretical Analysis
We consider both intra-view and inter-view scenarios.
Case 1 (Intra-view Validity:): We prove this by contradiction. Suppose honest participants $p_{i}$ and $p_{j}$ commit different proposals propprop and $prop^{\prime}$ respectively with the same sequence number $sq$ in view $v$. For participant $p_{i}$ to commit propprop, it must have received $2f+1$ matching Commit messages for some digest $\delta=h(prop)$ (i.e., $committed\text{-}local(\delta,v,sq)=true$). Similarly, for participant $p_{j}$ to commit $prop^{\prime}$, it must have received $2f+1$ matching Commit messages for some digest $\delta^{\prime}=h(prop^{\prime})$ (i.e., $committed\text{-}local(\delta^{\prime},v,sq)=true$). Since there are at most $f$ Byzantine participants, By the quorum intersection property (2(2f+1)n=(4f+2)(3f+1)=f+12(2f+1)-n=(4f+2)-(3f+1)=f+1), there exists at least one honest participant $p^{*}$ that sent both Commit messages for $\delta$ and $\delta^{\prime}$. However, honest participants only send one Commit message per sequence number per view, leading to a contradiction.
@@ VIII-A1 Experimental Settings
Datasets and Models. We evaluate the performance of the proposed ACuMPA scheme on three widely used visual recognition benchmark datasets: MNIST [28], Fashion-MNIST [29], and CIFAR-10 [30].
@@
For model evaluation, we employ three popular neural network architectures: a Convolutional Neural Network (CNN) [31], ResNet [32], and AlexNet [33]. Detailed specifications of the model structures are presented in Table II.